OpenStudy (anonymous):

g(f(x))=|sinx| , f(g(x))=(sin (square root(x))whole square, than find f(x) and g(x).

OpenStudy (anonymous):

\[g(f(x))=|sinx| , f(g(x))= (\sin \sqrt{x})^{2}\]

OpenStudy (anonymous):

equations are as shown below

OpenStudy (anonymous):

above, not below

OpenStudy (anonymous):

sorry missfont

OpenStudy (owen3):

Do you know that $$ \sqrt{u^2}= \left| u\right| $$

OpenStudy (owen3):

That's a big hint

OpenStudy (anonymous):



OpenStudy (anonymous):

yes i know

OpenStudy (owen3):

Well? I'm not allowed to just give you the answer.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so tell me the concept

OpenStudy (owen3):

Look at the hint above.

OpenStudy (anonymous):

yes i saw that square root x = mod x

OpenStudy (anonymous):

so, after this